OpenStudy (anonymous):

there are 14 boys and 20 girls...whats the probability of picking a boy

OpenStudy (anonymous):

14 over 34 7 over 17

OpenStudy (mathstudent55):

\(P(event) = \dfrac{number ~of~ways~~desired~outcome~can ~happen}{total~number~of ~outcomes} \) The total number of outcomes is 14 + 20 = 34 (you cna pick any of 34 boys and girls) The number of desired outcomes is 14 (only 14 of the picks are boys) \(P(picking~a~boy) = \dfrac{14}{34} \) You can reduce the fraction.

OpenStudy (anonymous):

ya the fraction reduced would be 7 over 17